# Optimal Execution in Intraday Energy Markets under Hawkes Processes with Transient Impact

### v1

@@ 2.1 The Marked Hawkes Price Model
Figure 4: Sample paths of the intensity process in (6) and the associated counting process $N_{t}$ with $\alpha=1$, $\beta=5$, and $\lambda_{0}=0.7$.
Using an exponential decay kernel $\phi^{ij}(t)=\alpha\mathop{}\!\textnormal{e}^{-\beta t}$ for $t\geq 0$ and $\beta\geq 0$, the conditional intensities are:
$\bm{\lambda}_{t}$ $=\bar{\bm{\lambda}}_{t}\!+\alpha\int_{0}^{t}\!\mathop{}\!\textnormal{e}^{-\beta(t-u)}\begin{pmatrix}\varphi_{s}(Y_{u})&\varphi_{c}(Y_{u})\\ \varphi_{c}(Y_{u})&\varphi_{s}(Y_{u})\end{pmatrix}\!\!\begin{pmatrix}\mathop{}\!\textnormal{d}N^{(1)}_{u}\\ \mathop{}\!\textnormal{d}N^{(2)}_{u}\end{pmatrix}$ (7)
where $\bm{\lambda}_{t}=(\lambda_{t}^{(1)},\lambda_{t}^{(2)})^{\top}$ is the bivariate intensity process, $\bar{\bm{\lambda}}_{t}=(\lambda_{\infty}^{(1)}(t),\lambda_{\infty}^{(2)}(t))^{\top}$ denotes the time-dependent baseline intensities with $\lambda^{(i)}_{\infty}(t)\geq 0$, $m_{1}=\int_{K}y\,\nu(dy)$ is the expected mark size, and $\varphi_{s},\varphi_{c}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ are measurable, positive functions satisfying:
$i_{\cdot}$ $:=\int_{K}\varphi^{(\cdot)}\left(\dfrac{y}{m_{1}}\right)\nu(dy)<\infty.$ (8)
@@
with superscripts $(s)$ and $(c)$ denoting self-excitation and cross-excitation functions, respectively. The self-excitation function $\varphi^{(s)}$ captures the effect of past events of the same type (buy or sell) on the current intensity of that type, while the cross-excitation function $\varphi^{(c)}$ captures the effect of past events of the opposite type on the current intensity. The conditions in (8) ensure that the expected contributions from the excitation functions are finite, which is necessary for the process to be well-defined. For stationarity of the bivariate Hawkes process, it is required that $i_{s}+i_{c}<\beta$, see [6], ensuring that the intensities remain finite over time.
@@
To model the volatility components of the price model, we further assume that the second moments of the excitation functions are finite:
$\int_{K}\left[\varphi^{(\cdot)}\left(\dfrac{y}{m_{1}}\right)\right]^{2}\nu(dy)<\infty,\quad$
The parameters $i_{s}$ and $i_{c}$ represent the average self-excitation and cross-excitation effects, respectively.
Assuming we know the shape of the intraday seasonality, we consider the deseasonalized intensity process $(\tilde{\lambda}^{+}_{m,t},\tilde{\lambda}^{-}_{m,t})$, obtained by removing the baseline intensity to focus on the excitation contributions:
$\tilde{\bm{\lambda}}_{t}$ $=\alpha\int_{0}^{t}\mathop{}\!\textnormal{e}^{-\beta(t-u)}\begin{pmatrix}\varphi_{s}(Y_{u})&\varphi_{c}(Y_{u})\\ \varphi_{c}(Y_{u})&\varphi_{s}(Y_{u})\end{pmatrix}\begin{pmatrix}dN^{(1)}_{u}\\ dN^{(2)}_{u}\end{pmatrix}$ (9)
where $\tilde{\bm{\lambda}}_{t}=\bm{\lambda}_{t}-\bar{\bm{\lambda}}_{t}$ denotes the deseasonalized intensity process, with $\bm{\lambda}_{t}=(\lambda^{(1)}_{t},\lambda^{(2)}_{t})^{\top}$ and $\bar{\bm{\lambda}}_{t}=(\lambda^{(1)}_{\infty}(t),\lambda^{(2)}_{\infty}(t))^{\top}$ being the raw and baseline intensities respectively. The functions $\varphi_{s},\varphi_{c}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ are measurable and positive, and $m_{1}=\int_{K}y\,\nu(dy)$ denotes the expected mark size.
@@
We recall the empirical observation of increasing intensity and volatility as time to maturity decreases, known as the Samuelson effect (see Figure 2). Incorporating a time-varying baseline intensity in (7) effectively captures this phenomenon. While most studies use a parametric approach, typically assuming an exponential function for the baseline intensity, we propose a more flexible non-parametric approach using spline functions, such as cosine bump functions.
@@ 2.2 Optimal Execution
In this section, we describe the general framework of the optimal execution problem without specifying the exact dynamics of the underlying processes. We consider a single asset, specifically an hourly product in the intraday energy market, and denote its price at time $t$ by $P_{t}$. We assume that $P_{t}$ can be decomposed into a fundamental price component $S_{t}$ and a transient price deviation $D_{t}$, such that:
@@
Following the framework proposed by Alfonsi et al. [2], let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. We define $N_{t}$ as the cumulative signed volume of past market orders excluding the trader’s own trades up to time $t$, given by $N_{t}=N_{t}^{+}-N_{t}^{-}$, where $N_{t}^{+}$ and $N_{t}^{-}$ are the cumulative volumes of buy and sell orders from other market participants, respectively. By convention, a buy order increases $N_{t}$, while a sell order decreases it. We assume that $N_{t}$ is a càdlàg process.
@@
where $\epsilon\in[0,1]$ represents the proportion of the trader’s price impact that is permanent. We may assume $\epsilon=\mu$, implying that the trader’s impact mirrors that of other market participants.
Analogously to some existing literature (see [4, 31]; for intraday markets, see [26]), we will assume a throughout a block-shaped limit order book. When the trader executes an order of size $\nu\in\mathbb{R}$ (with $\nu>0$ for a buy order and $\nu<0$ for a sell order), they incur a cost:
[…]
@@
Considering the cost associated with the trader’s liquidation strategy over the time interval $[0,T]$, we derive the cost functional for a block-shaped limit order book as:
[…]
where $[X]_{u}$ denotes the quadratic variation of $X$ up to time $u$. We refer to $C(X)$ as the pathwise cost associated with a given strategy $X$.
@@
As we can see from the definition of the market order arrival rates given by Equation (9), the trader’s orders (i.e., our trading activity) do not impact the jump rates $\lambda^{+}$ and $\lambda^{-}$, since there is no additional $dX_{t}$ term influencing the arrival rates of subsequent market orders. This contrasts with the market orders issued by other traders (the market activity which we calibrate directly from the LOB). The main reason for this modeling choice is tractability and the assumption that each individual trader tends to send several orders of the same sign in a row, which in turn creates autocorrelation in the signs of trades [2] and this effect is stronger than the mutual excitation between different traders. However, recent work by Horst et al. [15] accounts for the fact that the trading activity of a large producer, whose trades may impact the market, triggers child orders and endogenously affects the future order flow.
@@
In this setting, we study the optimal execution problem under the assumption that buy and sell market order flows follow Hawkes processes. The arrival times of buy and sell orders are modeled as the jump times of two dependent Hawkes processes, $N^{+}$ (buy) and $N^{-}$ (sell), with conditional instantaneous intensities defined by (9). The dependence between these processes is captured by the impact matrix:
$H:=\begin{pmatrix}\varphi^{(s)}(y/m_{1})&\varphi^{(c)}(y/m_{1})\\ \varphi^{(c)}(y/m_{1})&\varphi^{(s)}(y/m_{1})\end{pmatrix},$ (17)
where $\varphi^{(s)}$ and $\varphi^{(c)}$ are the self-excitation and cross-excitation functions, respectively, and $m_{1}=\int_{K}y\,\nu(dy)$ is the mean mark. This matrix parameterizes the excitation effects, accounting for the dependence between the arrival of buy and sell market orders.
@@
We derive the optimal execution strategy in the most general setting. Later, in the empirical results and backtesting section, we will explore different modeling approaches. Notably, the optimal execution problem can be solved explicitly within the Marked Hawkes price model due to its linear price impact and the exponential decay kernel of the Hawkes processes, which ensure Markovian dynamics. Furthermore, for completely monotone decay kernels, Markovian dynamics can also be recovered, as shown by Alfonsi et al. [4]. The derivation is carried out in the deseasonalized case, where the baseline intensity is constant.
@@
Before introducing the optimal trading strategy and the main theorem of this section, we first introduce some auxiliary variables that will be useful later on. Instead of working directly with $\lambda_{t}^{+}$ and $\lambda_{t}^{-}$, we define the processes:
$\delta_{t}=\lambda_{t}^{+}-\lambda_{t}^{-},\quad\gamma_{t}=\lambda_{t}^{+}+\lambda_{t}^{-}.$
@@
The process $\delta$ describes the order flow imbalance and $\gamma$ is the total order intensity. These processes satisfy the following stochastic differential equations:
$d\delta_{t}$ $=-\beta\delta_{t}\,dt+dI_{t},$
$d\gamma_{t}$ $=-\beta\left(\gamma_{t}-2\lambda_{\infty}\right)dt+d\bar{I}_{t},$
@@
where $\lambda_{\infty}=\lambda_{\infty}^{+}+\lambda_{\infty}^{-}$ is the sum of the baseline intensities, and the processes $I_{t}$ and $\bar{I}_{t}$ are given by:
$I_{t}$ $=\int_{0}^{t}\left[\left(\phi_{s}-\phi_{c}\right)\frac{dN_{s}^{+}}{m_{1}}-\left(\phi_{s}-\phi_{c}\right)\frac{dN_{s}^{-}}{m_{1}}\right],$ (18)
$\bar{I}_{t}$ $=\int_{0}^{t}\left[\left(\phi_{s}+\phi_{c}\right)\frac{dN_{s}^{+}}{m_{1}}+\left(\phi_{s}+\phi_{c}\right)\frac{dN_{s}^{-}}{m_{1}}\right],$ (19)
with $m_{1}=\int_{K}y\,\nu(dy)$ being the mean mark size, and $\phi_{s}$, $\phi_{c}$ are constants representing the self-excitation and cross-excitation effects, respectively.
@@
Let $\left(\tau_{i}\right)_{i\in\mathbb{N}}$ denote the ordered random jump times of the counting process $N$, with $\tau_{0}=0$. For $t\in[0,T]$, let $\chi_{t}$ denote the total number of jumps of $I$ that have occurred between time 00 and $t$. Then, the process $\delta_{t}$ can be expressed as:
@@
$\delta_{t}=\delta_{0}\mathop{}\!\textnormal{e}^{-\beta t}+\sum_{\ell=1}^{\chi_{t}}\mathop{}\!\textnormal{e}^{-\beta(t-\tau_{\ell})}\Delta I_{\tau_{\ell}}=\mathop{}\!\textnormal{e}^{-\beta t}\left(\delta_{0}+\Theta_{\chi_{t}}\right),$ (20)
@@
with $\Theta_{0}=0$. This representation shows that for $i\geq 0$ and $t\in[\tau_{i},\tau_{i+1})$, the quantity $\mathop{}\!\textnormal{e}^{\beta t}\delta_{t}=\delta_{0}+\Theta_{i}$ depends only on the number of jumps up to time $t$, i.e., $\chi_{t}=i$.
The jump intensity of $(N_{t})_{t\geq 0}$ is determined by the Markov processes $\delta_{t}$ and $\gamma_{t}$, taking values in $\mathbb{R}$ and $\times\mathbb{R}+$, respectively, as defined in Equations (18). The state variables of the optimization problem are then $\left(X_{t},D_{t},S_{t},\delta_{t},\gamma_{t}\right)$, and the control is the trading strategy $\left(X_{t}\right){t\in[0,T]}$, with $X_{0}=x_{0}$.
@@
$\mathcal{C}(t,Z_{t})=\inf_{X\in\mathcal{A}_{t}}\mathbb{E}\left[C(t,X)\,\big{|}\,Z_{t}=(x,d,z,\delta,\gamma)\right],$ (23)
for $Z_{t}=(X_{t},D_{t},S_{t},\delta_{t},\gamma_{t})$. The terminal condition for the value function is:
$\mathcal{C}(T,x,d,z,\delta,\gamma)=-(d+z)x+\dfrac{1}{2}x^{2}.$ (24)
The optimal strategy $X^{*}$ can be derived analytically in this model due to the Markovian and affine structure of the state space and the fact that the cost function (objective) is a linear-quadratic function of the state variables [3]. We also refer the reader to the work presented in [34] for a general analysis of affine point processes.
A continuously differentiable function $\mathcal{C}(t,x,d,z,\delta,\gamma)$ is determined a priori along with an admissible strategy $X^{*}$. By defining the process $\Pi_{t}(X)$ as
[…]
we verify the following important criterion: $\Pi_{t}(X)$ is a submartingale for any admissible strategy $X$, and $\Pi_{t}(X^{*})$ is a martingale where $X^{*}$ is the solution of the optimal control problem (23).
@@
$\zeta(y)$ =$=\begin{cases}1,&y=0,\\ \frac{1-\mathop{}\!\textnormal{e}^{-y}}{y},&y\neq 0,\end{cases}\qquad\omega(y)$ =$=\begin{cases}\frac{1}{2},&y=0,\\ \frac{\mathop{}\!\textnormal{e}^{-y}-1+y}{y^{2}},&y\neq 0.\end{cases}$ (26)
Let $h:=T-t$. Then, the optimal strategy satisfies
$X_{t}^{*}$ $=-(1-\mu)^{-1}[1+\rho h]D_{t}+\frac{m_{1}}{2\rho}[2+\rho h]$
×$\quad\times\delta_{t}\left[1+\frac{\rho h}{2+\rho h}\left(\zeta\left(h\eta\right)+\mu\rho h\,\omega\left(h\eta\right)\right)\right],$
where $\delta_{t}=\lambda_{t}^{+}-\lambda_{t}^{-}$ is the intensity order flow imbalance as before.

### v2

@@ 2.1 The Marked Hawkes Price Model
Figure 3: Sample paths of the intensity process and the associated counting process $N_{t}$ with $\alpha=1$, $\beta=5$, and $\lambda_{\infty}=0.7$. Sample paths of a univariate Hawkes process $N_{t}$ with intensity $\lambda_{t}=\lambda_{\infty}+\int_{0}^{t}\alpha\mathop{}\!\textnormal{e}^{-\beta(t-s)}\,dN_{s}$ for $\alpha=1$, $\beta=5$, and $\lambda_{0}=0.7$.
Note that the family of counting processes $\{N_{m}(t)\}_{m=1}^{M}$ forms a $2M$-dimensional Hawkes process. In our setting, however, we abstract from cross-product interactions and do not model correlations across different delivery products. Hence, no cross-excitation effects are included between products. For ease of notation, we therefore fix a single delivery product and drop the index $m$. Buy and sell arrivals are then modeled by a bivariate Hawkes process with exponential decay kernels. Specifically, we let
$\bm{\lambda}_{t}:=\begin{pmatrix}\lambda_{t}^{+}\\ \lambda_{t}^{-}\end{pmatrix},\qquad\bar{\bm{\lambda}}_{t}:=\begin{pmatrix}\lambda_{\infty}^{+}(t)\\ \lambda_{\infty}^{-}(t)\end{pmatrix}$,
denote the conditional and baseline intensities (with the latter possibly time-varying). We assume that the kernel matrix is separable, of the form
$\phi^{\pm}(u,y)=\phi(u)\odot\Phi(y),$
where $\odot$ denotes the Hadamard product, $\varphi(u)$ captures the exponential decay in time, and $\Phi(y)$ encodes the mark dependence via
$\Phi(y)=\begin{pmatrix}\varphi^{(s)}(y)&\varphi^{(c)}(y)\\ \varphi^{(c)}(y)&\varphi^{(s)}(y)\end{pmatrix},\quad\varphi^{(s)}(y),\varphi^{(c)}(y)\geq 0.$ (4)
Here, $\varphi^{(s)}$ and $\varphi^{(c)}:K\to\mathbb{R}_{+}$ denote the mark-dependent self- and cross-excitation amplitudes, where the superscripts $(s)$ and $(c)$ refer to self- and cross-excitation, respectively. In the special case where $\varphi^{(s)}$ and $\varphi^{(c)}$ are constant functions, the model reduces to the standard bivariate Hawkes process. Throughout, we assume the following moment condition to hold
$i_{\cdot}:=\int_{K}\varphi^{(\cdot)}\!\left(\frac{y}{m_{1}}\right)\nu(dy)<\infty,$ (5)
where $m_{1}=\int_{K}y\,\nu(dy)$ denotes the expected mark size. In the remainder, we standardize the mark sizes by $m_{1}$ and let the conditional intensity vector be given by
$\bm{\lambda}_{t}=\bar{\bm{\lambda}}_{t}+\int_{0}^{t}\phi(t-u)\,\odot\,\Phi\begin{pmatrix}\tfrac{dN^{+}(u)}{m_{1}}\\ \tfrac{dN^{-}(u)}{m_{1}}\end{pmatrix}$, (6)
where the integral is understood component-wise and we note that $\int_{0}^{t}\varphi^{(s)}\left(\tfrac{dN^{+}(u)}{m_{1}}\right)=\int_{0}^{t}\int_{K}\varphi^{(s)}\left(y/m_{1}\right)\pi^{+}\left(du,dy\right)$.
@@
The self-excitation function $\varphi^{(s)}$ captures the effect of past events of the same type (buy or sell) on the current intensity of that type, while the cross-excitation function $\varphi^{(c)}$ captures the effect of past events of the opposite type on the current intensity. The conditions in (5) ensure that the expected contributions from the excitation functions are finite, which is necessary for the process to be well-defined.
@@
To model the volatility components of the price model, we further assume that the second moments of the excitation functions are finite, i.e.
$\int_{K}\left[\varphi^{(\cdot)}\left(\dfrac{y}{m_{1}}\right)\right]^{2}\nu(dy)<\infty.$
Assuming the shape of intraday seasonality is known, we define the deseasonalized intensity process as
$\tilde{\bm{\lambda}}_{t}:=\bm{\lambda}_{t}-\bar{\bm{\lambda}}_{t},\qquad t\geq 0,$ (7)
where $\tilde{\bm{\lambda}}_{t}=(\tilde{\lambda}^{+}_{t},\tilde{\lambda}^{-}_{t})^{\top}$ represents the residual intensity obtained by removing the baseline $\bar{\bm{\lambda}}_{t}$.
@@
We recall the empirical observation of increasing intensity and volatility as time to maturity decreases, known as the Samuelson effect (see Figure 2(a)). Incorporating a time-varying baseline intensity in (6) effectively captures this phenomenon. While most studies use a parametric approach, typically assuming an exponential function for the baseline intensity, we propose a more flexible non-parametric approach using spline functions, such as cosine bump functions, see Section 3.1 below.
@@ 2.2 Optimal Execution
In this section, we describe the framework of the optimal execution problem in our model class. Throughout this section, we consider a single asset, specifically an hourly product in the intraday energy market, and denote its price at time $t$ by $P_{t}$. We assume that $P_{t}$ can be decomposed into a fundamental price component $S_{t}$ given by (2) and a transient price deviation $D_{t}$, such that:
@@
We define $N_{t}$ as the cumulative signed volume of past market orders excluding the trader’s own trades up to time $t$, given by
$N_{t}:=N_{t}^{+}-N_{t}^{-},$
where $N_{t}^{+}$ and $N_{t}^{-}$ are the cumulative volumes of buy and sell orders from other market participants, respectively. By convention, a buy order increases $N_{t}$ and $S_{t}$, while a sell order decreases them.
@@
where $\epsilon\in[0,1]$ represents the proportion of the trader’s price impact that is permanent. For simplicity, and since we do not impose additional assumptions on the information structure of other market participants, we set $\epsilon=\mu$, i.e., the trader’s impact is assumed to be identical to that of the other participants.
Analogously to some existing literature (see [Alfonsi2012, Obizhaeva2005]; for intraday markets, see [KZ20]), we will assume throughout a block-shaped limit order book. When the trader executes an order of size $\nu\in\mathbb{R}$ (with $\nu>0$ for a buy order and $\nu<0$ for a sell order), they incur a cost:
$\pi_{t}(\nu)=\int_{0}^{\nu}\left(P_{t}+y\right)dy=P_{t}\nu+\frac{1}{2}\nu^{2},$ (13)
@@
Let $D_{X}$ be the (at most countable) set of jump times of $X$ and $\Delta X_{\tau}:=X_{\tau_{+}}-X_{\tau}$. Summing the block costs over trades before terminal time $T$ and adding the terminal block needed to liquidate any residual inventory $X_{T}$ at time $T$ yields the following liquidation cost
$C(X)$ $=\int_{[0,T)}P_{u}\,dX_{u}+\frac{1}{2}\sum_{\tau\in D_{X}\cap[0,T)}(\Delta X_{\tau})^{2}$
$\quad-P_{T}X_{T}+\frac{1}{2}X_{T}^{2}.$ (14)
$=\int_{0}^{T}\!P_{u}\,dX_{u}+\!\frac{1}{2}\int_{0}^{T}\!d[X]_{u}\!-P_{T}X_{T}\!+\frac{1}{2}X_{T}^{2}.$
For a simple strategy with jumps $\{\tau_{i}<T\}$, the total paid is $\sum_{i}\big(P_{\tau_{i}}\Delta X_{\tau_{i}}+\tfrac{1}{2}(\Delta X_{\tau_{i}})^{2}\big)$ which follows by (13) and block-shaped assumption. If $X_{T}\neq 0$, liquidation at $T$ adds the terminal execution cost $P_{T}(-X_{T})+\tfrac{1}{2}X_{T}^{2}$. Passing to the càdlàg limit gives (2.2). The Stieltjes integral $\int_{[0,T)}P_{u}\,dX_{u}$ already accounts for the $P_{\tau_{i}}\Delta X_{\tau_{i}}$ terms. Hence by denoting with $X]_{t}$ the quadratic variation of $X$ up to time $t$ and letting $X_{t}=X_{t}^{c}+\sum_{\tau\in D_{X}\cap[0,T)}\left(X_{\tau}^{+}-X_{\tau}\right)$ the last equality in (2.2) follows.
@@
As we can see from the definition of the market order arrival rates given by equation (7), the trader’s orders (i.e., our trading activity) do not impact the jump rates $\lambda^{+}$ and $\lambda^{-}$, since there is no additional $dX_{t}$ term influencing the arrival rates of subsequent market orders. This contrasts with the market orders issued by other traders (the market activity which we calibrate directly from the LOB). The main reason for this modeling choice is tractability and the assumption that each individual trader tends to send several orders of the same sign in a row, which in turn creates autocorrelation in the signs of trades [Alfonsi2016] and this effect is stronger than the mutual excitation between different traders. However, recent work by Horst et al. [CHTH23] accounts for the fact that the trading activity of a large market participant, whose trades may impact the market, triggers child orders and endogenously affects the future order flow.
@@
In this setting, we study the optimal execution problem under the assumption that buy and sell market order flows follow Hawkes processes. The arrival times of buy and sell orders are modeled as the jump times of two dependent Hawkes processes, $N^{+}$ (buy) and $N^{-}$ (sell), with conditional instantaneous intensities defined by (7). The dependence between these processes is captured by the impact matrix in Equation (4).
@@
We now derive the optimal execution strategy in the most general setting. In the empirical results and backtesting section, we will study two different modeling approaches. A key feature of the Marked Hawkes price model is that the optimal execution problem can be solved in closed form: the linear price impact and the exponential decay kernels of the Hawkes processes ensure Markovian dynamics. More generally, Markovian representations can also be recovered for completely monotone decay kernels, as shown by Alfonsi et al. [Alfonsi2012]. For clarity throughout this section, we work in the deseasonalized case assuming the same constant baseline intensities for buy and sell side.
@@
Before introducing the optimal trading strategy and the main theorem, we define auxiliary variables that simplify the dynamics. Instead of working directly with $\lambda_{t}^{+}$ and $\lambda_{t}^{-}$, we consider
$\kappa_{t}=\lambda_{t}^{+}-\lambda_{t}^{-},\qquad\gamma_{t}=\lambda_{t}^{+}+\lambda_{t}^{-},$
@@
where $\kappa_{t}$ describes the order-flow imbalance and $\gamma_{t}$ the total order intensity. By the variation-of-constant formula, these processes satisfy the following stochastic differential equations:
$d\kappa_{t}$ $=-\beta\kappa_{t}\,dt+dI_{t},$
$d\gamma_{t}$ $=-\beta\big(\gamma_{t}-\lambda_{\infty}\big)dt+d\bar{I}_{t},$
@@
where $\lambda_{\infty}=\lambda_{\infty}^{+}+\lambda_{\infty}^{-}$ is the sum of baseline intensities with $\lambda_{\infty}^{+}=\lambda_{\infty}^{-}$, and the processes $I_{t}$ and $\bar{I}_{t}$ encode the impact of arrivals and are given by
$I_{t}$ $=\!\int_{0}^{t}\!\big((\varphi^{(s)}-\varphi^{(c)})\tfrac{dN_{s}^{+}}{m_{1}}-(\varphi^{(s)}-\varphi^{(c)})\tfrac{dN_{s}^{-}}{m_{1}}\big),$ (15)
$\bar{I}_{t}$ $=\!\int_{0}^{t}\!\big((\varphi^{(s)}+\varphi^{(c)})\tfrac{dN_{s}^{+}}{m_{1}}+(\varphi^{(s)}+\varphi^{(c)})\tfrac{dN_{s}^{-}}{m_{1}}\big).$ (16)
@@
Let $\left(\tau_{i}\right)_{i\in\mathbb{N}}$ denote the ordered random jump times of the counting process $N$, with $\tau_{0}=0$. For $t\in[0,T]$, let $\chi_{t}$ denote the total number of jumps of $I$ that have occurred between time 0 and $t$. Then, the process $\kappa_{t}$ can be expressed as:
@@
$\kappa_{t}=\kappa_{0}\mathop{}\!\textnormal{e}^{-\beta t}+\sum_{\ell=1}^{\chi_{t}}\mathop{}\!\textnormal{e}^{-\beta(t-\tau_{\ell})}\Delta I_{\tau_{\ell}}=\mathop{}\!\textnormal{e}^{-\beta t}\left(\kappa_{0}+\Theta_{\chi_{t}}\right),$ (17)
@@
with $\Theta_{0}=0$. This representation shows that for $i\geq 0$ and $t\in[\tau_{i},\tau_{i+1})$, the quantity $\mathop{}\!\textnormal{e}^{\beta t}\kappa_{t}=\kappa_{0}+\Theta_{i}$ depends only on the number of jumps up to time $t$, i.e., $\chi_{t}=i$.
The jump intensity of $(N_{t})_{t\geq 0}$ is determined by the Markov processes $\kappa_{t}$ and $\gamma_{t}$ as defined in Equations (15) taking values in $\mathbb{R}$ and $\mathbb{R}_{+}$, respectively. The state variables of the optimization problem are then $\left(X_{t},D_{t},S_{t},\kappa_{t},\gamma_{t}\right)$, and the control is the trading strategy $\left(X_{t}\right)_{t\in[0,T]}$, with $X_{0}=x_{0}$.
@@
$\mathcal{C}(t,Z_{t})=\inf_{X\in\mathcal{A}_{t}}\mathbb{E}\left[C(t,X)\,\big|\,Z_{t}=(x,d,z,\kappa,\gamma)\right],$ (20)
for $Z_{t}=(X_{t},D_{t},S_{t},\kappa_{t},\gamma_{t})$. The terminal condition for the value function is:
$\mathcal{C}(T,x,d,z,\kappa,\gamma)=-(d+z)x+\dfrac{1}{2}x^{2}.$ (21)
The optimal strategy $X^{*}$ can be derived analytically in this model due to the Markovian and affine structure of the state space and the fact that the cost function (objective) is a linear-quadratic function of the state variables [AlfonsiExtension]. For this, one defines a continuously differentiable function $\mathcal{C}(t,x,d,z,\kappa,\gamma)$ a priori along with an admissible strategy $X^{*}$, and define the process $\Pi_{t}(X)$ as
$\!\!\Pi_{t}(X)=\!\int_{0}^{t}\!P_{u}\,dX_{u}\!+\!\frac{1}{2}\int_{0}^{t}\!d[X,X]_{u}+\mathcal{C}(t,Z_{t}).$ (22)
Finally, one has to verify that $\Pi_{t}(X)$ is a submartingale for any admissible strategy $X$, and $\Pi_{t}(X^{*})$ is a martingale for $X^{*}$ the solution of the optimal control problem (20). This optimality criterion gives then rise to the following theorem.
@@
[…]
Let h:=Tth:=T-t. Then, the optimal strategy satisfies
$X_{t}^{*}$ $=-(1-\mu)^{-1}[1+\rho h]D_{t}+\frac{m_{1}}{2\rho}[2+\rho h]$ (24)
$\quad\times\kappa_{t}\left[1+\frac{\rho h}{2+\rho h}\left(\zeta\left(h\eta\right)+\mu\rho h\,\omega\left(h\eta\right)\right)\right],$
where $\kappa_{t}=\lambda_{t}^{+}-\lambda_{t}^{-}$ is the intensity order flow imbalance as before.
Proof.
This follows from [Alfonsi2016, Theorem 4.1] and [AlfonsiExtension, Theorem 2.2] under the assumption of mono-exponential kernels for the Hawkes intensity and the propagator kernel. ∎